Dada la fórmula I = prt, si p = 8000, r = 0.6 y t = 1, ¿cuál es el valor de "I"?
Answer by Boreal(15235)

I=prt=8000*0.6*1=$4800.
pienso que r es 0.06 o 6%.
Entonces I=8000*0.06*1=$480